OpenStudy (anonymous):

Yuri solved the equation below incorrectly. Explain the step where he went wrong. fifteen times b minus one equals thirty Step 1: fifteen times b minus one minus one equals thirty minus one Step 2: fifteen times b equals twenty-nine Step 3: fifteen times b divided by fifteen equals twenty-nine divided by fifteen Step 4: b equals twenty-nine fifteenths

OpenStudy (cwrw238):

15b - 1 = 30 15b - 1 - 1 = 30-1 15b - 2 = 29 step 1 : 1 should be added to both sides in order to get 15b = 31

OpenStudy (cwrw238):

the error is in step 2 - taking 1 from each side does not give 15b = 29